OpenStudy (anonymous):

express -2 - 2i in trigonoetric form

OpenStudy (crashonce):

@majestymmm is 'i' an imaginary number?

OpenStudy (tkhunny):

Where is (-2,-2) on an x-y coordinate plane?

OpenStudy (bradely):

-2-2i a+bi magnitude,r =sqrt((-2)^2+(-2)^2) =2sqrt(2) \[\theta=\pi +\tan ^{-1}(-2/-2)\] since a<0 =pi+pi/4 -5pi/4 Trigonometric form ,r =2sqrt(2)(cos5pi/4+isin5pi/4) Source: http://www.mathskey.com/question2answer/